a. If the interest rate is 5.4% per year, approximately how long will it take for your money to quadruple in value? (Use theWhats the correct answer to part b?
Answer #1

a.

4 = 1(1.054)t

Using Log,

t = log(4)/log(1.054)

t = 26.36 years

b.

Increases;

Real Rate = 0.054 - 0.037 = 1.7%

Value = 1(1.017)26.36

Value = 1.5595

Increase = 55.95%